OpenStudy (anonymous):

the square of a number is equal to 20 more than the number. find all such numbers

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Let the number be x then: \(x^2=20+x\), implies that \(x^2-x-20=0\). This can be factored into \((x-5)(x+4)=0\). Hence \(5\) and \(-4\) are the only two numbers that satisfy the given conditions.

OpenStudy (dumbcow):

n^2 = n+20 quadratic equation set equal to 0 n^2 -n -20 = 0 factor (n-5)(n+4)=0

OpenStudy (anonymous):

thankss =]